# Jian-hao Kanga, Zhun Goub and Nan-jing Huangc

### v1

@@ 6 Conclusions
Since $Y^{\gamma}(t,x,m_{1})=-\frac{1}{\gamma}e^{g_{1}^{\gamma}(t)(x+\beta m_{1})+g_{2}^{\gamma}(t)}$, where $\gamma\in[\epsilon_{1},\infty)$, $g_{1}^{\gamma}(t)=-\gamma e^{(A+\beta)(T-t)}$ and $g_{2}^{\gamma}(t)$ is given by the equation (4.26), we can derive that $Y^{\gamma}\in C^{1,2,1}(\mathcal{D})$ and then
$\left|Y^{\gamma}(t,X^{\hat{u}}(t),M_{1}^{\hat{u}}(t))\right|^{2}=$ $\left|\frac{1}{\gamma^{2}}e^{2g_{1}^{\gamma}(t)\left(X^{\hat{u}}(t)+\beta M_{1}^{\hat{u}}(t)\right)+2g_{2}^{\gamma}(t)}\right|$
$\leq$ $Ke^{2g_{1}^{\gamma}(t)\left(X^{\hat{u}}(t)+\beta M_{1}^{\hat{u}}(t)\right)}$
$=$ $Ke^{2g_{1}^{\gamma}(t)\left(x_{0}e^{At}+\int^{t}_{0}e^{A(t-s)}\left[X^{\hat{u}}(s)\hat{\pi}(s)(\mu-r)+a\eta+a\eta_{2}\hat{q}(s)+BM_{1}^{\hat{u}}(s)+CM_{2}^{\hat{u}}(s)\right]\mathrm{d}s\right)}$
$\times e^{2g_{1}^{\gamma}(t)\int^{t}_{0}e^{A(t-s)}\left[b\hat{q}(s)\mathrm{d}W_{1}(s)+\hat{\pi}(s)X^{\hat{u}}(s)\sigma\mathrm{d}W_{2}(s)\right]+2\beta g_{1}^{\gamma}(t)M_{1}^{\hat{u}}(t)}$
$\leq$ $Ke^{-2\gamma\int^{t}_{0}e^{(A+\beta)T-\beta t-As}\left(BM_{1}^{\hat{u}}(s)+CM_{2}^{\hat{u}}(s)\right)\mathrm{d}s-2\gamma\int^{t}_{0}e^{-\beta(t-s)}\left[b\bar{q}\mathrm{d}W_{1}(s)+\bar{\pi}\sigma\mathrm{d}W_{2}(s)\right]}$
$\times e^{-2\beta\gamma e^{(A+\beta)(T-t)}M_{1}^{\hat{u}}(t)},$
where we employ the boundedness of $g_{2}^{\gamma}(t)$, $X^{\hat{u}}(t)\hat{\pi}(t)$ and $\hat{q}(t)$, $K$ is a positive constant that can take different values in different lines and will be used thereafter when there is no ambiguity, $\bar{q}=\frac{a\eta_{2}}{b^{2}\mathbb{E}[\gamma]}$ and $\bar{\pi}=\frac{\mu-r}{\sigma^{2}\mathbb{E}[\gamma]}$. Let $M_{3}(t)=e^{-16\gamma\int^{t}_{0}e^{(A+\beta)T-\beta t-As}BM_{1}^{\hat{u}}(s)\mathrm{d}s}$ and $M_{4}(t)=e^{-16\gamma\int^{t}_{0}e^{(A+\beta)T-\beta t-As}CM_{2}^{\hat{u}}(s)\mathrm{d}s}$. Then we have
$\mathrm{d}M_{3}(t)=-16\gamma Be^{(A+\beta)T-\beta t}\left[e^{-At}M_{1}^{\hat{u}}(t)-\beta\int^{t}_{0}e^{-As}M_{1}^{\hat{u}}(s)\mathrm{d}s\right]M_{3}(t)\mathrm{d}t$
with $M_{3}(0)=1$. By Lemma 2.1 in [32] and Lemma 3.1, we know that there exists a unique solution $M_{3}(\cdot)\in\mathbb{S}^{2}_{\mathcal{F}}(0,T;\mathbb{R})$ to the above equation for $M_{3}(t)$. Similarly, we can obtain that $M_{4}(\cdot)\in\mathbb{S}^{2}_{\mathcal{F}}(0,T;\mathbb{R})$. Therefore, similar to the analysis in Lemma 3.1, we can also obtain that $M_{3}(\cdot),M_{4}(\cdot)\in\mathbb{S}^{1}_{\mathcal{F}}(0,T;\mathbb{R})$. Besides, letting $M_{5}(t)=e^{-8\beta\gamma e^{(A+\beta)(T-t)}M_{1}^{\hat{u}}(t)}$, by the equation (2.3), one can arrive at
$\mathrm{d}M_{5}(t)=-8\beta\gamma e^{(A+\beta)(T-t)}\left[X^{\hat{u}}(t)-(A+\beta+\alpha)M_{1}^{\hat{u}}(t)-e^{-\alpha h}M_{2}^{\hat{u}}(t)\right]M_{5}(t)\mathrm{d}t$
with $M_{5}(0)=e^{-8\beta\gamma e^{(A+\beta)T}\frac{x_{0}(1-e^{-\alpha h})}{\alpha}}$. Reapplying Lemma 2.1 in [32] and Lemma 3.1, we can conclude that the above equation for $M_{5}(t)$ admits a unique solution $M_{5}(\cdot)\in\mathbb{S}^{2}_{\mathcal{F}}(0,T;\mathbb{R})$ and thus $M_{5}(\cdot)\in\mathbb{S}^{1}_{\mathcal{F}}(0,T;\mathbb{R})$.
Furthermore, denoting $M_{6}(t)=-4\gamma\int^{t}_{0}e^{\beta s}b\bar{q}\mathrm{d}W_{1}(s)$, $M_{7}(t)=-4\gamma\int^{t}_{0}e^{\beta s}\bar{\pi}\sigma\mathrm{d}W_{2}(s)$ and $M_{8}(t)=e^{e^{-\beta t}(M_{6}(t)+M_{7}(t))}$, we can derive that
$\mathbb{E}\left[\sup_{t\in[0,T]}\left|M_{8}(t)\right|\right]\leq\mathbb{E}\left[e^{\sup\limits_{t\in[0,T]}e^{-\beta t}\times\sup\limits_{t\in[0,T]}(M_{6}(t)+M_{7}(t))}\right]\leq\mathbb{E}\left[\sup_{t\in[0,T]}e^{M_{6}(t)+M_{7}(t)}\right],$
if $\sup\limits_{t\in[0,T]}(M_{6}(t)+M_{7}(t))>0$. On the other hand, if $\sup\limits_{t\in[0,T]}(M_{6}(t)+M_{7}(t))<0$, then one has
$\sup_{t\in[0,T]}[e^{-\beta t}(M_{6}(t)+M_{7}(t))]\leq e^{-\beta T}\sup\limits_{t\in[0,T]}(M_{6}(t)+M_{7}(t))$
and thus $\mathbb{E}\left[\sup\limits_{t\in[0,T]}\left|M_{8}(t)\right|\right]\leq e^{e^{-\beta T}}\mathbb{E}\left[\sup\limits_{t\in[0,T]}e^{M_{6}(t)+M_{7}(t)}\right]$. Let $M_{9}(t)=e^{M_{6}(t)+M_{7}(t)}$. Then, we know that
$\mathrm{d}M_{9}(t)=M_{9}(t)\left[8\gamma^{2}e^{2\beta t}\left(b^{2}\bar{q}^{2}+\bar{\pi}^{2}\sigma^{2}\right)\mathrm{d}t-4\gamma e^{\beta t}\left(b\bar{q}\mathrm{d}W_{1}(s)+\bar{\pi}\sigma\mathrm{d}W_{2}(s)\right)\right]$
with $M_{9}(0)=1$. It follows from Theorem 1.3.15 in [29] that there exists a unique strong solution $M_{9}(\cdot)\in\mathbb{S}^{2}_{\mathcal{F}}(0,T;\mathbb{R})$ to the above equation. Ulteriorly, we have $M_{9}(\cdot)\in\mathbb{S}^{1}_{\mathcal{F}}(0,T;\mathbb{R})$, which implies that
$\mathbb{E}\left[\sup_{t\in[0,T]}\left|M_{8}(t)\right|\right]<\infty$
for the cases where $\sup\limits_{t\in[0,T]}(M_{6}(t)+M_{7}(t))>0$ and $\sup\limits_{t\in[0,T]}(M_{6}(t)+M_{7}(t))<0$. If $\sup\limits_{t\in[0,T]}(M_{6}(t)+M_{7}(t))=0$, then the conclusion also holds trivially. In addition, one has
$\mathbb{E}\left[\sup_{t\in[0,T]}\left|Y^{\gamma}(t,X^{\hat{u}}(t),M_{1}^{\hat{u}}(t))\right|^{2}\right]$
@@
$\;\;\;\left.\left.\times e^{-2\gamma\int^{t}_{0}e^{-\beta(t-s)}\left[b\bar{q}\mathrm{d}W_{1}(s)+\bar{\pi}\sigma\mathrm{d}W_{2}(s)\right]-2\beta\gamma e^{(A+\beta)(T-t)}M_{1}^{\hat{u}}(t)}\right\}\right]$
$\leq$ $K\mathbb{E}\left[\sup_{t\in[0,T]}\left(M_{3}(t)+M_{4}(t)+M_{5}(t)+M_{8}(t)\right)\right]$
$<$ $\infty,$
where we use the fact that $M_{3}(\cdot),M_{4}(\cdot),M_{5}(\cdot)\in\mathbb{S}^{1}_{\mathcal{F}}(0,T;\mathbb{R})$. We now proceed to obtain that
@@
$=$ $\mathbb{E}\left[\int^{T}_{0}\left|-\frac{1}{\gamma}e^{g_{1}^{\gamma}(t)\left(X^{\hat{u}}(t)+\beta M_{1}^{\hat{u}}(t)\right)+g_{2}^{\gamma}(t)}\left[\frac{\partial g_{1}^{\gamma}(t)}{\partial t}\left(X^{\hat{u}}(t)+\beta M_{1}^{\hat{u}}(t)\right)+\frac{\partial g_{2}^{\gamma}(t)}{\partial t}\right]\right|\mathrm{d}t\right]$
@@
in which the inequality holds because $\frac{\partial g_{1}^{\gamma}(t)}{\partial t}$ and $\frac{\partial g_{2}^{\gamma}(t)}{\partial t}$ are deterministic and bounded. Since $X^{\hat{u}}(\cdot)\in\mathbb{S}^{2}_{\mathcal{F}}(0,T;\mathbb{R})$, we have
@@
$=$ $\left(\mathbb{E}\left[\sup_{t\in[0,T]}\left|Y^{\gamma}(t,X^{\hat{u}}(t),M_{1}^{\hat{u}}(t))\right|^{2}\right]\times\mathbb{E}\left[\sup_{t\in[0,T]}\left|X^{\hat{u}}(t)\right|^{2}\right]\right)^{\frac{1}{2}}$
$<$ $\infty.$
Similarly, by using $M_{1}^{\hat{u}}(\cdot)\in\mathbb{S}^{2}_{\mathcal{F}}(0,T;\mathbb{R})$, we can also obtain
@@
By equations (C) and (4.27) and $X^{\hat{u}}(\cdot),M_{1}^{\hat{u}}(\cdot)\in\mathbb{S}^{1}_{\mathcal{F}}(0,T;\mathbb{R})$, we have $U,H\in C^{1,2,1}(\mathcal{D})$ and
@@
Appendix F F Proof of Theorem 4.2
@@
$\left\{\begin{aligned} \mathrm{d}X^{\hat{u}}(t)=&\left[AX^{\hat{u}}(t)+BM_{1}^{\hat{u}}(t)+CM_{2}^{\hat{u}}(t)+\left(\frac{(\mu-r)^{2}}{\sigma^{2}\varpi(t,\Gamma)}+\frac{a^{2}\eta^{2}_{2}}{b^{2}\varpi(t,\Gamma)}\right)(X^{\hat{u}}(t)+\beta M_{1}^{\hat{u}}(t))\right]\mathrm{d}t\\ &+\frac{a\eta_{2}}{b\varpi(t,\Gamma)}(X^{\hat{u}}(t)+\beta M_{1}^{\hat{u}}(t))\mathrm{d}W_{1}(t)+\frac{\mu-r}{\sigma\varpi(t,\Gamma)}(X^{\hat{u}}(t)+\beta M_{1}^{\hat{u}}(t))\mathrm{d}W_{2}(t),\;t\in[\tau,T],\\ X^{\hat{u}}(t)=&\psi(t-\tau),\;t\in[\tau-h,\tau].\end{aligned}\right$. (F.57)
By Proposition 2.1 in [28], one can obtain that the SDDE (F.57) admits a unique strong solution $X^{\hat{u}}(\cdot)\in\mathbb{S}^{p}_{\mathcal{F}}(0,T;\mathbb{R})$ for $p\geq 2$. This indicates that the condition (i) in Definition 3.1 holds. In light of $X^{\hat{u}}(\cdot)\in\mathbb{S}^{2}_{\mathcal{F}}(0,T;\mathbb{R})$ and Lemma 3.1, one can also get that $M_{1}^{\hat{u}}(\cdot),\;M_{2}^{\hat{u}}(\cdot)\in\mathbb{S}^{2}_{\mathcal{F}}(0,T;\mathbb{R})$ and $X^{\hat{u}}(\cdot),\;M_{1}^{\hat{u}}(\cdot),\;M_{2}^{\hat{u}}(\cdot)\in\mathbb{S}^{1}_{\mathcal{F}}(0,T;\mathbb{R})$. Moreover, by using equations (2.3) and (F.57), one has
@@
$+\left(X^{\hat{u}}(t)+\beta M_{1}^{\hat{u}}(t)\right)\left[M_{10}(t)\mathrm{d}t+M_{11}(t)\mathrm{d}W_{1}(t)+M_{12}(t)\mathrm{d}W_{2}(t)\right],$ (F.58)
where $M_{10}(t)=\frac{(\mu-r)^{2}}{\sigma^{2}\varpi(t,\Gamma)}+\frac{a^{2}\eta^{2}_{2}}{b^{2}\varpi(t,\Gamma)}$, $M_{11}(t)=\frac{a\eta_{2}}{b\varpi(t,\Gamma)}$ and $M_{12}(t)=\frac{\mu-r}{\sigma\varpi(t,\Gamma)}$. Applying the condition $(A2)$, the equation (F) can be rewritten as
$\mathrm{d}\left[X^{\hat{u}}(t)+\beta M_{1}^{\hat{u}}(t)\right]=\left(X^{\hat{u}}(t)+\beta M_{1}^{\hat{u}}(t)\right)\left[\left(A+\beta+M_{10}(t)\right)\mathrm{d}t+M_{11}(t)\mathrm{d}W_{1}(t)+M_{12}(t)\mathrm{d}W_{2}(t)\right].$
@@
$X^{\hat{u}}(t)+\beta M_{1}^{\hat{u}}(t)=$ $\left(x_{0}+\frac{\beta x_{0}(1-e^{-\alpha h})}{\alpha}\right)\exp\left\{\int^{t}_{0}\left(A+\beta+M_{10}(s)-\frac{(M_{11}(s))^{2}+(M_{12}(s))^{2}}{2}\right)\mathrm{d}s\right\}$
$\times\exp\left\{\int^{t}_{0}\left(M_{11}(s)\mathrm{d}W_{1}(s)+M_{12}(s)\mathrm{d}W_{2}(s)\right)\right\},$
which shows that $X^{\hat{u}}(t)+\beta M_{1}^{\hat{u}}(t)>0$ for $t\in[0,T]$ since $x_{0}$, $\beta$, $\alpha$ and $h$ are positive. Thus, it follows from the equation (4.43) that $\hat{q}(t)>0$ and $\{(\hat{q}(t),\hat{\pi}(t))\}_{t\in[0,T]}$ is $\mathcal{F}_{t}$-progressively measurable and continuous. In addition, since $X^{\hat{u}}(\cdot),\;M_{1}^{\hat{u}}(\cdot)\in\mathbb{S}^{2}_{\mathcal{F}}(0,T;\mathbb{R})$, one has
@@
On the other hand, we find that
$|M_{1}^{\hat{u}}(t)|=\left|\int^{0}_{-h}e^{\alpha s}X^{\hat{u}}(t+s)\mathrm{d}s\right|\leq K\max_{s\in[-h,0]}|X^{\hat{u}}(t+s)|\leq K\|X^{\hat{u}}_{t}\|$
and $|M_{2}^{\hat{u}}(t)|=|X^{\hat{u}}(t-h)|\leq\max\limits_{s\in[-h,0]}|X^{\hat{u}}(t+s)|=\|X^{\hat{u}}_{t}\|$. Clearly, $|X^{\hat{u}}(t)|\leq\|X^{\hat{u}}_{t}\|$. Because $Y^{\gamma_{i}}(t,x,m_{1})=\frac{1}{1-\gamma_{i}}(g^{\gamma_{i}}(t))^{\gamma_{i}}(x+\beta m_{1})^{1-\gamma_{i}}$, where $g^{\gamma_{i}}(t)$ is captured by the equation (4.41) and bounded on $[0,T]$, we can derive that $Y^{\gamma_{i}}\in C^{1,2,1}(\mathcal{D})$ and
@@
Hence, $Y^{\gamma_{i}}_{t}(t,X^{\hat{u}}(t),M_{1}^{\hat{u}}(t))\in\mathbb{L}^{1}_{\mathcal{F}}(0,T;\mathbb{R})$. Additionally, letting $M_{13}(t)=AX^{\hat{u}}(t)+BM_{1}^{\hat{u}}(t)+CM_{2}^{\hat{u}}(t)$ and $M_{14}(t)=X^{\hat{u}}(t)-\alpha M_{1}^{\hat{u}}(t)-e^{-\alpha h}M_{2}^{\hat{u}}(t)$, by the condition $(A2)$ and applying a similar proof procedure as for $Y^{\gamma_{i}}_{t}(t,X^{\hat{u}}(t),M_{1}^{\hat{u}}(t))\in\mathbb{L}^{1}_{\mathcal{F}}(0,T;\mathbb{R})$, one can get that
$\mathbb{E}\left[\int^{T}_{0}\left|M_{13}(t)Y^{\gamma_{i}}_{x}(t,X^{\hat{u}}(t),M_{1}^{\hat{u}}(t))+M_{14}(t)Y^{\gamma_{i}}_{m_{1}}(t,X^{\hat{u}}(t),M_{1}^{\hat{u}}(t))\right|\mathrm{d}t\right]$
$=$ $\mathbb{E}\left[\int^{T}_{0}\left|M_{13}(t)(g^{\gamma_{i}}(t))^{\gamma_{i}}\left(X^{\hat{u}}(t)+\beta M_{1}^{\hat{u}}(t)\right)^{-\gamma_{i}}+M_{14}(t)\beta(g^{\gamma_{i}}(t))^{\gamma_{i}}\left(X^{\hat{u}}(t)+\beta M_{1}^{\hat{u}}(t)\right)^{-\gamma_{i}}\right|\mathrm{d}t\right]$
@@
which means that $M_{13}(t)Y^{\gamma_{i}}_{x}(t,X^{\hat{u}}(t),M_{1}^{\hat{u}}(t))+M_{14}(t)Y^{\gamma_{i}}_{m_{1}}(t,X^{\hat{u}}(t),M_{1}^{\hat{u}}(t))\in\mathbb{L}^{1}_{\mathcal{F}}(0,T;\mathbb{R})$. Analogously, it can be shown that $M_{15}(t)Y^{\gamma_{i}}_{x}(t,X^{\hat{u}}(t),M_{1}^{\hat{u}}(t))+M_{16}(t)Y^{\gamma_{i}}_{xx}(t,X^{\hat{u}}(t),M_{1}^{\hat{u}}(t))\in\mathbb{L}^{1}_{\mathcal{F}}(0,T;\mathbb{R})$ with $M_{15}(t)=\hat{\pi}(t)(\mu-r)X^{\hat{u}}(t)+a\eta_{2}\hat{q}(t)$ and $M_{16}(t)=0.5\left(b^{2}\hat{q}^{2}(t)+\hat{\pi}^{2}(t)(X^{\hat{u}}(t))^{2}\sigma^{2}\right)$. Moreover, by using a similar proof process as for $Y^{\gamma_{i}}_{t}(t,X^{\hat{u}}(t),M_{1}^{\hat{u}}(t))\in\mathbb{L}^{1}_{\mathcal{F}}(0,T;\mathbb{R})$, one can arrive at

### v2

@@ 7 Conclusions
$\mathrm{d}\left[X^{\hat{u}}(t)+\beta M_{1}^{\hat{u}}(t)\right]=$ $\left[(A+\beta)(X^{\hat{u}}(t)+\beta M_{1}^{\hat{u}}(t))+\frac{(\mu-r)^{2}}{\sigma^{2}\mathbb{E}[\gamma]}e^{-(A+\beta)(T-t)}+a\eta+\frac{a^{2}\eta^{2}_{2}}{b^{2}\mathbb{E}[\gamma]}e^{-(A+\beta)(T-t)}\right]\mathrm{d}t$
$+\frac{a\eta_{2}}{b\mathbb{E}[\gamma]}e^{-(A+\beta)(T-t)}\mathrm{d}W_{1}(t)+\frac{\mu-r}{\sigma\mathbb{E}[\gamma]}e^{-(A+\beta)(T-t)}\mathrm{d}W_{2}(t).$
Then, based on the above SDDE, we can obtain that
$\left|Y^{\gamma}(t,X^{\hat{u}}(t),M_{1}^{\hat{u}}(t))\right|^{4}=$ $\left|\frac{1}{\gamma^{4}}e^{4g_{1}^{\gamma}(t)\left(X^{\hat{u}}(t)+\beta M_{1}^{\hat{u}}(t)\right)+4g_{2}^{\gamma}(t)}\right|$
$\leq$ $Ke^{4g_{1}^{\gamma}(t)\left(X^{\hat{u}}(t)+\beta M_{1}^{\hat{u}}(t)\right)}$
$=$ $Ke^{4g_{1}^{\gamma}(t)\left((x_{0}+\beta m_{10})e^{(A+\beta)t}+\int^{t}_{0}e^{(A+\beta)(t-s)}\left[X^{\hat{u}}(s)\hat{\pi}(s)(\mu-r)+a\eta+a\eta_{2}\hat{q}(s)\right]\mathrm{d}s\right)}$
$\times e^{4g_{1}^{\gamma}(t)\int^{t}_{0}e^{(A+\beta)(t-s)}\left[b\hat{q}(s)\mathrm{d}W_{1}(s)+\hat{\pi}(s)X^{\hat{u}}(s)\sigma\mathrm{d}W_{2}(s)\right]}$
$\leq$ $Ke^{-4\gamma\int^{t}_{0}\left[b\bar{q}\mathrm{d}W_{1}(s)+\bar{\pi}\sigma\mathrm{d}W_{2}(s)\right]},$
where we employ the boundedness of $g_{2}^{\gamma}(t)$, $X^{\hat{u}}(t)\hat{\pi}(t)$ and $\hat{q}(t)$, $\bar{q}=\frac{a\eta_{2}}{b^{2}\mathbb{E}[\gamma]}$ and $\bar{\pi}=\frac{\mu-r}{\sigma^{2}\mathbb{E}[\gamma]}$. Define $\widetilde{M}_{4}(t)=e^{-4\gamma\int^{t}_{0}\left[b\bar{q}\mathrm{d}W_{1}(s)+\bar{\pi}\sigma\mathrm{d}W_{2}(s)\right]}$. Because $\bar{q}$ and $\bar{\pi}$ are constants and $W_{1}$ and $W_{2}$ are independent, we find that
$\widetilde{M}_{4}(T)=\underbrace{e^{8\gamma^{2}\int^{T}_{0}\left(b^{2}\bar{q}^{2}+\bar{\pi}^{2}\sigma^{2}\right)\mathrm{d}s}}_{constant}\times\underbrace{e^{-8\gamma^{2}\int^{T}_{0}\left(b^{2}\bar{q}^{2}+\bar{\pi}^{2}\sigma^{2}\right)\mathrm{d}s-4\gamma\int^{T}_{0}\left[b\bar{q}\mathrm{d}W_{1}(s)+\bar{\pi}\sigma\mathrm{d}W_{2}(s)\right]}}_{martingale}.$
Hence, $\mathbb{E}\left[\widetilde{M}_{4}(T)\right]<\infty$. By Burkhölder-Davis-Gundy inequality, we know that
$\mathbb{E}\left[\sup_{t\in[0,T]}\left|Y^{\gamma}(t,X^{\hat{u}}(t),M_{1}^{\hat{u}}(t))\right|^{4}\right]$ $\leq K\mathbb{E}\left[\left|Y^{\gamma}(T,X^{\hat{u}}(T),M_{1}^{\hat{u}}(T))\right|^{4}\right]\leq K\mathbb{E}\left[\widetilde{M}_{4}(T)\right]<\infty.$ (F.72)
Proceeding as in the proof of Lemma 3.1, one arrives at $Y^{\gamma}(\cdot,X^{\hat{u}}(\cdot),M_{1}^{\hat{u}}(\cdot))\in\mathbb{S}^{2}_{\mathcal{F}}(0,T;\mathbb{R})$.
[…]
@@
$=$ $\mathbb{E}\left[\int^{T}_{0}\left|Y^{\gamma}(t,X^{\hat{u}}(t),M_{1}^{\hat{u}}(t))\left[\frac{\partial g_{1}^{\gamma}(t)}{\partial t}\left(X^{\hat{u}}(t)+\beta M_{1}^{\hat{u}}(t)\right)+\frac{\partial g_{2}^{\gamma}(t)}{\partial t}\right]\right|\mathrm{d}t\right]$
@@
in which the inequality holds because $\frac{\partial g_{1}^{\gamma}(t)}{\partial t}$ and $\frac{\partial g_{2}^{\gamma}(t)}{\partial t}$ are deterministic and bounded. Since $Y^{\gamma}(\cdot,X^{\hat{u}}(\cdot),M_{1}^{\hat{u}}(\cdot))\in\mathbb{S}^{2}_{\mathcal{F}}(0,T;\mathbb{R})$ and $X^{\hat{u}}(\cdot)\in\mathbb{S}^{2}_{\mathcal{F}}(0,T;\mathbb{R})$, we have
@@
$=$ $\left(\mathbb{E}\left[\sup_{t\in[0,T]}\left|Y^{\gamma}(t,X^{\hat{u}}(t),M_{1}^{\hat{u}}(t))\right|^{2}\right]\times\mathbb{E}\left[\sup_{t\in[0,T]}\left|X^{\hat{u}}(t)\right|^{2}\right]\right)^{\frac{1}{2}}<\infty.$
Because $X^{\hat{u}}(\cdot)\in\mathbb{S}^{2}_{\mathcal{F}}(0,T;\mathbb{R})$, one has $M_{1}^{\hat{u}}(\cdot)\in\mathbb{S}^{2}_{\mathcal{F}}(0,T;\mathbb{R})$ by the proof process of Theorem 4.1. Similarly, we can also obtain
@@
[…]
@@
Appendix G G Proof of Theorem 5.2
@@
$\left\{\begin{aligned} \mathrm{d}X^{\hat{u}}(t)=&\left[AX^{\hat{u}}(t)+BM_{1}^{\hat{u}}(t)+CM_{2}^{\hat{u}}(t)+\left(\frac{(\mu-r)^{2}}{\sigma^{2}\varpi(t,\Gamma)}+\frac{a^{2}\eta^{2}_{2}}{b^{2}\varpi(t,\Gamma)}\right)(X^{\hat{u}}(t)+\beta M_{1}^{\hat{u}}(t))\right]\mathrm{d}t\\ &+\frac{a\eta_{2}}{b\varpi(t,\Gamma)}(X^{\hat{u}}(t)+\beta M_{1}^{\hat{u}}(t))\mathrm{d}W_{1}(t)+\frac{\mu-r}{\sigma\varpi(t,\Gamma)}(X^{\hat{u}}(t)+\beta M_{1}^{\hat{u}}(t))\mathrm{d}W_{2}(t),\;t\in[\tau,T],\\ X^{\hat{u}}(t)=&\psi(t-\tau),\;t\in[\tau-h,\tau].\end{aligned}\right$. (G.73)
By Proposition 2.1 in [28], one can obtain that the SDDE (G.73) admits a unique strong solution $X^{\hat{u}}(\cdot)\in\mathbb{S}^{p}_{\mathcal{F}}(0,T;\mathbb{R})$ for $p\geq 2$. This indicates that the condition (i) in Definition 3.1 holds. Moreover, by using equations (2.3) and (G.73), one has
@@
$+\left(X^{\hat{u}}(t)+\beta M_{1}^{\hat{u}}(t)\right)\left[\widetilde{M}_{5}(t)\mathrm{d}t+\widetilde{M}_{6}(t)\mathrm{d}W_{1}(t)+\widetilde{M}_{7}(t)\mathrm{d}W_{2}(t)\right],$ (G.74)
where $\widetilde{M}_{5}(t)=\frac{(\mu-r)^{2}}{\sigma^{2}\varpi(t,\Gamma)}+\frac{a^{2}\eta^{2}_{2}}{b^{2}\varpi(t,\Gamma)}$, $\widetilde{M}_{6}(t)=\frac{a\eta_{2}}{b\varpi(t,\Gamma)}$ and $\widetilde{M}_{7}(t)=\frac{\mu-r}{\sigma\varpi(t,\Gamma)}$. Applying the condition $(A2)$, the equation (G) can be rewritten as
$\mathrm{d}\left[X^{\hat{u}}(t)+\beta M_{1}^{\hat{u}}(t)\right]=\left(X^{\hat{u}}(t)+\beta M_{1}^{\hat{u}}(t)\right)\left[\left(A+\beta+\widetilde{M}_{5}(t)\right)\mathrm{d}t+\widetilde{M}_{6}(t)\mathrm{d}W_{1}(t)+\widetilde{M}_{7}(t)\mathrm{d}W_{2}(t)\right].$
@@
$X^{\hat{u}}(t)+\beta M_{1}^{\hat{u}}(t)=$ $\left(x_{0}+\frac{\beta x_{0}(1-e^{-\alpha h})}{\alpha}\right)\exp\left\{\int^{t}_{0}\left(A+\beta+\widetilde{M}_{5}(s)-\frac{(\widetilde{M}_{6}(s))^{2}+(\widetilde{M}_{7}(s))^{2}}{2}\right)\mathrm{d}s\right\}$
$\times\exp\left\{\int^{t}_{0}\left(\widetilde{M}_{6}(s)\mathrm{d}W_{1}(s)+\widetilde{M}_{7}(s)\mathrm{d}W_{2}(s)\right)\right\},$
which shows that $X^{\hat{u}}(t)+\beta M_{1}^{\hat{u}}(t)>0$ for $t\in[0,T]$ since $x_{0}$, $\beta$, $\alpha$ and $h$ are positive. Thus, it follows from the equation (5.53) that $\hat{q}(t)>0$ and $\{(\hat{q}(t),\hat{\pi}(t))\}_{t\in[0,T]}$ is $\mathcal{F}_{t}$-progressively measurable and continuous. In addition, since $X^{\hat{u}}(\cdot)\in\mathbb{S}^{2}_{\mathcal{F}}(0,T;\mathbb{R})$, one can derive that $M_{1}^{\hat{u}}(\cdot)\in\mathbb{S}^{2}_{\mathcal{F}}(0,T;\mathbb{R})$ by the proof process of Theorem 4.1. Therefore, one has
@@
[…]
@@
Hence, $Y^{\gamma_{i}}_{t}(t,X^{\hat{u}}(t),M_{1}^{\hat{u}}(t))\in\mathbb{L}^{1}_{\mathcal{F}}(0,T;\mathbb{R})$. Additionally, letting $\widetilde{M}_{8}(t)=AX^{\hat{u}}(t)+BM_{1}^{\hat{u}}(t)+CM_{2}^{\hat{u}}(t)$ and $\widetilde{M}_{9}(t)=X^{\hat{u}}(t)-\alpha M_{1}^{\hat{u}}(t)-e^{-\alpha h}M_{2}^{\hat{u}}(t)$, by the condition $(A2)$ and applying a similar proof procedure as for $Y^{\gamma_{i}}_{t}(t,X^{\hat{u}}(t),M_{1}^{\hat{u}}(t))\in\mathbb{L}^{1}_{\mathcal{F}}(0,T;\mathbb{R})$, one can get that
$\mathbb{E}\left[\int^{T}_{0}\left|\widetilde{M}_{8}(t)Y^{\gamma_{i}}_{x}(t,X^{\hat{u}}(t),M_{1}^{\hat{u}}(t))+\widetilde{M}_{9}(t)Y^{\gamma_{i}}_{m_{1}}(t,X^{\hat{u}}(t),M_{1}^{\hat{u}}(t))\right|\mathrm{d}t\right]$
$=$ $\mathbb{E}\left[\int^{T}_{0}\left|\widetilde{M}_{8}(t)(g^{\gamma_{i}}(t))^{\gamma_{i}}\left(X^{\hat{u}}(t)+\beta M_{1}^{\hat{u}}(t)\right)^{-\gamma_{i}}+\widetilde{M}_{9}(t)\beta(g^{\gamma_{i}}(t))^{\gamma_{i}}\left(X^{\hat{u}}(t)+\beta M_{1}^{\hat{u}}(t)\right)^{-\gamma_{i}}\right|\mathrm{d}t\right]$
@@
which means that $\widetilde{M}_{8}(t)Y^{\gamma_{i}}_{x}(t,X^{\hat{u}}(t),M_{1}^{\hat{u}}(t))+\widetilde{M}_{9}(t)Y^{\gamma_{i}}_{m_{1}}(t,X^{\hat{u}}(t),M_{1}^{\hat{u}}(t))\in\mathbb{L}^{1}_{\mathcal{F}}(0,T;\mathbb{R})$. Analogously, it can be shown that $\widetilde{M}_{10}(t)Y^{\gamma_{i}}_{x}(t,X^{\hat{u}}(t),M_{1}^{\hat{u}}(t))+\widetilde{M}_{11}(t)Y^{\gamma_{i}}_{xx}(t,X^{\hat{u}}(t),M_{1}^{\hat{u}}(t))\in\mathbb{L}^{1}_{\mathcal{F}}(0,T;\mathbb{R})$ with $\widetilde{M}_{10}(t)=\hat{\pi}(t)(\mu-r)X^{\hat{u}}(t)+a\eta_{2}\hat{q}(t)$ and $\widetilde{M}_{11}(t)=0.5\left(b^{2}\hat{q}^{2}(t)+\hat{\pi}^{2}(t)(X^{\hat{u}}(t))^{2}\sigma^{2}\right)$. Moreover, by using a similar proof process as for $Y^{\gamma_{i}}_{t}(t,X^{\hat{u}}(t),M_{1}^{\hat{u}}(t))\in\mathbb{L}^{1}_{\mathcal{F}}(0,T;\mathbb{R})$, one can arrive at